Circular No. 8128

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ec IN NGC 5990 A.-C. Chen, S.-Y. Wu, W.-H. Sun, Y.-W. Cheng, S.-Y. Wu, W.-H. Sun,

Y.-W. Cheng, M.-T. Chang, H.-C. Lin, and C.-C. Liu, National Central University, Chung-li, Taiwan, report the discovery of an apparent supernova (V = 16.4) on B, V, and R CCD images obtained on May 2.801 and 3.662 UT with the Lulin Observatory 1-m telescope. SN 2003ec is located at α = 15^{h}46^{m}16^{s}.67, δ = +2^{o}24′48′′.7 (equinox 2000.0), which is 4′′.2 east and 7′′.2 south of the center of NGC 5990. Nothing appears at this location on a 1982 Digitized Sky Survey image.

(3782) CELLE W. H. Ryan, New Mexico Institute of Mining and Technology and

New Mexico Highlands University, reports that photometric observations of (3782) Celle were obtained during 2001 Sept. 18.3–19.5, 2002 Dec. 9.3– 11.4, 2003 Jan. 4.1–6.2, and Jan. 26.1–Feb. 1.3 UT at the 1.8-m Vatican Advanced Technology Telescope as part of a larger survey by a collaboration that also includes E. V. Ryan, C. T. Martinez, and L. Stewart. Analysis of these data reveals a normal rotational lightcurve (period 3.84 hr, amplitude 0.10–0.15 mag) superimposed onto deeper attenuation events lasting for ∼ 2.5–3.5 hr that varied in amplitude from 0.15–0.3 mag. The onset of these events was observed on 2003 Jan. 5.25, 27.34, 28.10, 30.39, and 31.15. An event in progress was also observed on 2002 Dec. 9.4. The attenuations were of two distinct types that can clearly be identified as primary and secondary occultations/eclipses, similar to those that have been previously observed in known minor-planet binary systems (Pravec et al. 2000, Icarus 146, 190). In particular, the 2003 Jan. 5.25 event was indicative of a complete occultation of the secondary component of such a binary system. The data are interpreted as clear evidence that (3782) is actually an asynchronous binary system with a primary-to-secondary diameter ratio of 0.42 ± 0.02 and an orbital period of 36.57 ± 0.03 hr. Since (3782) has been associated with the Vesta family both dynamically (Zappalá et al. 1984, A.J. 107,

772) and spectroscopically (Bus and Binzel 2002, Icarus 158, 146), this is the first identified binary amongst the so-called ‘Vesta chips’.

COMET 156P/RUSSELL-LINEAR Comet P/2000 QD_{181} = 2000 XV_{43} = 1986 R1 = 1993 WU (Russell-

LINEAR), announced on IAUC 8118, has been given the permanent number 156P (MPC 48317).

2003 May 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green